In a multiple- choice test, each question has five options. Students will get 3 points for each correct answer. lose 1 point for each incorrect answer, and receive no points for unanswered questions. A student does not know the correct answer for one question. Is it to her advantage or disadvantage to guess an answer?
1) Show calculations for expected value.
2) Show all work to solve the problem.

Answer by stanbon(75887)

In a multiple- choice test, each question has five options. Students will get 3 points for each correct answer. lose 1 point for each incorrect answer, and receive no points for unanswered questions. A student does not know the correct answer for one question. Is it to her advantage or disadvantage to guess an answer?
1) Show calculations for expected value.
Random Point Values: 3 , -1 , 0
Corresponding Probs: 1/5,4/5, 0
----------------------------------------
E(x) = (1/5)*3 + (4/5)*-1 + 0 = (3/5)-(4/5) = -1/5 point per question answered.
--------------
Cheers,
Stan H.
==============

2) Show all work to solve the problem.
done